the fundamental note produce by an open organ pipe has a frequency n. the fundamental note produces by a closed organ pipe of same length will be.

Grade:12th pass

1 Answers

Arun
6 years ago
Dear student
for open organ pipe fundamental frequency is f1 = v/2l.
for closed organ pipe fundamental frequency is f2 = v/4l.
f1/f2 = 2
Hence
Frequency of closed organ pipes = n/2
Regards
Arun (askIITians forum expert)